# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ Introduction
This is the idea of the aforementioned monotonicity-duality method initiated by Harvey and Lawson in [23] in the constant coefficient pure second order case. It has developed into a robust and general method which can also be transported to every nonlinear degenerate elliptic PDE which is compatibile with a given potential theory. The reader is invited to consult [24, 9, 10, 8, 11] for many of these developments. We will derive this method from first principles and then discuss the constant coefficient pure second order and gradient-free cases, which also includes the characterizzation of the dual cone subharmonics, which are the subaffine and subaffine-plus functions, respectively. The second setting is when there is the minimal monotonicity that every subsequation possesses and one obtains a weaker form of the comparison principle, called strict comparison. In the variable coefficient case, when the subharmonics are quasi-convex the proof of strict comparison is a simple consequence of the Summand Theorem. Two approaches to strict comparison for semicontinuous subharmonics are presented in the constant coefficient case. Both approaches use quasi-convex approximations; one implicitly and one explicitly.
This completes the synopsis of Parts I and II, where we note that there are also two appendices which, together, give a self-contained proof of Alexandrov’s Theorem. The proof depends on two ingredients: 1) second order differentiability at “non-critical points” of functions $f$ which are quadratic perturbations of a bounded convex function $u$, where the non-critical ponts are the non-critical values of the subdifferential of the convex Legendre transformation of $f$ (see Lemma A.7) and 2) a Lipschitz version of Sard’s theorem (see Lemma A.9), which shows that the set of “critical points” has Lebesgue measure zero.
We conclude this introduction with two considerations, which were previously mentioned. The first concerns parallel developments in nonlinear potential theories and nonlinear PDEs and the second concerns what is “new” in this primer, with respect to the two unpublished manuscripts [25] and [26].
As described above, the “hard analysis part” in nonlinear potential theories is contained in two deep results on quasi-convex functions; namely, Alexandov’s Theorem and the Jensen–Słodkowski Theorem. Indeed, Alexandov’s Theorem (and basic properties of subequations) implies the validity of the AE Theorem and hence the Subharmonic Addition Theorem. The Jensen–Słodkowski Theorem is used in both the proof of the hardest part of the Upper Contact Jet Theorem, partial upper semicontinuity of second derivatives, and the proof of the Summand Theorem. In addition, quasi-approximation via sup-convolution is a robust method for reducing a potential-theoretic question for semicontinuous functions to its quasi-convex counterpart. As is well known, the same basic ingredients play the same role in the conventional viscosity theory for nonlinear PDEs. Indeed, the proof of the crucial Theorem on Sums depends on Alexandrov’s Theorem, Jensen’s Lemma and “magical properties” of the sup-convolution.
A deep and impressive insight of Harvey and Lawson in [25] is that two seemingly different measure-theoretic results (implicit in the discussion above), when properly reformulated in the language of quadratic upper contact jets, are equivalent and can be merged into the aforementioned Jensen–Słodkowski Theorem 3.28. The first result due to Słodkowski [43], which was developed in the context of pluripotential theory, gives a lower bound on the lower Lebesgue density of sublevel sets of the generalized largest eigenvalue of the Hessian of a locally convex function (see Lemma 3.22). The second result due to Jensen [35], which was developed in the context of the uniqueness machinery for viscosity solutions of PDEs, says that near a strict local maximum of a quasi-convex function $w$ there is a set of positive Lebesgue measure of points where small linear perturbations of $w$ has a local maximum. Harvey and Lawson reformulate the Słodkowski and Jensen Lemmas in terms of quadratic upper contact jets, called here the HL-Słodkowski Lemma 3.21 and the HL-Jensen Lemma 3.23. These two reformulations are shown to be equivalent in Proposition 3.24, and are also equivalent to the full Jensen–Słodkowski Theorem, as noted in Remark 3.29. A lack of recognition of this equivalence represents a missed opportunity in the possible interplay between potential theories and PDEs. LABEL:pt1 offers a systematic study of these pivotal theorems and a presentation of some important related results.
@@
We conclude this introduction with some remarks on what is “new” in this primer with respect to the unpublished manuscripts [25] and [26]. In a general sense, almost everything one finds in the two preprints of Harvey and Lawson is included here, albeit sometimes arranged in a different manner, with modified, clarified and more detailed proofs. Many additional results and remarks have been included to enrich the presentation, which also takes into account the numerous developments in general nonlinear potential theory over the past decade.
@@
In a more technical sense, they are many things here that are not included in [25] and [26]. Some of these are well known, such as elements of convex function theory and the proof of the Lipschitz version of Sard’s theorem, which have been included in order to make the primer as self-contained as reasonably possible. Others are less well known, but needed for the exposition, such as the background material on subequations and their subharmonics in Sections LABEL:sec:conset, 6 and 7. Finally, there are some genuinely “new” things, many of which revolve around the following consideration.
@@
A key point in Harvey and Lawson’s study of Jensen–Słodkowski equivalences was to replace the spherical contact points considered by Słodkowski with paraboloidal contact points which are more natural for the viscosity theory. This “paraboloidal” machinery is examined in detail in Subsection 3.2. In addition, using this machinery, we augment the results in [25]. In particular, we formulate and prove a paraboloidal version of Słodkowski’s upper bound on the largest eigenvalue of the Hessian (Lemma 3.21) and we give a paraboloidal proof of Słodkowski’s density estimate (Lemma 3.1) and Słodkowski’s Largest Eigenvalue Theorem 3.5. Finally, in Subsection 3.6, making use of an important observation of Harvey [20], we prove that the general Jensen–Słodkowski Theorem 3.28 can be derived by Alexandrov’s maximum principle (Theorem 3.36), which is turn is derived from the Area formula for gradients of quasi-convex functions (Theorem 3.35). This leads to the interesting conclusion that, in the end, the hard analysis part rests on the validity of the area formula.
@@
Finally, an organic presentation of the quasi-convex approximation of $\mathcal{F}$ subharmonics is discussed in Section 8, which is then used to give an alternate proof of strict comparison per semicontinuous subharmonics (Theorem 9.17) by reducing to the quasi-convex version of strict comparison (Theorem 9.17).
PartI..
\@afterheading
Of course, this means that the restriction of $u$ to each segment $[x_{1},x_{2}]\subset X$ is a convex function of one variable. Moreover, the inequality (LABEL:Jensen_2) is a version of Jensen’s inequalty for convex functions.
@@
The proof of Lemma 1.3 is standard and typically done by induction on $N\geqslant 2$, where the case $N=2$ is (LABEL:Jensen_2).
Remark 1.4.
We also briefly remind the reader of the following obvious facts: affine functions are convex; nonnegative weighted sums of convex functions are convex; the (pointwise) supremum of any family of convex functions is convex.
@@
Geometrically, this means that the hyperplane which is the graph of the affine function $u(x)+\langle p,\cdot-x\rangle$ over $X$ is a supporting hyperplane from below at $x$ for the epigraph of $u$, for all $p\in\partial u(x)$. One can also say that if $u$ is subdifferentiable at $x$, then $x$ belongs to the (global) lower contact set of $u$. Moreover, one should note the simple but useful fact that $x$ is a global minimum point for $u$ if and only if $0\in\partial u(x)$. In particular, all functions are subdifferentiable at global minima.
@@
$\partial u\colon X\to\mathscr{P}({\mathbb{R}}^{n}),\quad x\mapsto\partial u(x)$
@@
$u\ \text{is convex}\ \ \Leftrightarrow\ \ \partial u(x)\neq\emptyset,\ \forall\,x\in X;$
@@
If $u$ is convex, then by the Hahn–Banach Theorem for any $x\in X$ there exists a supporting hyperplane for the convex set $C\vcentcolon=\overline{\operatorname{epi}}(u)$ at the point $(x,u(x))$; that is, there exist $(q,r)\in{\mathbb{R}}^{n}\times{\mathbb{R}}\setminus\{(0,0)\}$ such that
@@
which yields $-r^{-1}q\in\partial u(x)$, provided that $r\neq 0$. To show that in fact $r\neq 0$, notice that otherwise by inequality (1.4) one would have $\langle q,y-x\rangle\geqslant 0$ for each $y\in X$, which is impossible since $X$ is open. Conversely, if $x=\lambda x_{1}+(1-\lambda)x_{2}$ for some $\lambda\in[0,1]$, we have
@@
$u$ is bounded on every compact subset $K\subset X$, (1.5)
@@
$u$ is bounded on some neighborhood of every $x\in X$. (1.6)
@@
The equivalence of (1.5) and (1.6) is standard. Let $K\subset X$ be compact. By choosing any $x\in K$ and any $p\in\partial u(x)\neq\emptyset$, one has
$u(y)\geqslant u(x)+\langle p,y-x\rangle:=a_{x}(y),\ \ \text{for every}\ y\in X,$
where $a_{x}$ is affine and hence bounded from below on $K$ compact. Hence $u$ is bounded from below on $K$. For the local boundedness from above, let $x\in X$ be arbitrary and consider any compact cube $\mathcal{C}$ centered in $x$ such that $\mathcal{C}\subset X$. We denote by $\{v_{j}\}_{j=1}^{N}$ the collection of vertices of $\mathcal{C}$. Every closed cube is the (closed) convex hull of its vertices and hence for each $y\in\mathcal{C}$ one has
@@
$u(y)=u\Bigg{(}\sum_{j=1}^{N}t_{j}v_{j}\Bigg{)}\leqslant\sum_{j=1}^{N}t_{j}u(v_{j})\leqslant\max_{1\leqslant j\leqslant N}\lvert u(v_{j})\rvert<+\infty,$
@@
    the fiber $\partial u(x)\subset{\mathbb{R}}^{n}$ is closed and convex for each $x\in X$;
@@
$u(y)\geqslant u(x)+\langle p,y-x\rangle\quad\forall p\in\partial u(x),\ \forall\,y\in X$ (1.8)
@@
$u(x)\geqslant u(y)+\langle q,x-y\rangle\quad\forall q\in\partial u(y),\ \forall\,x\in X,$ (1.9)
@@
We will see that the chain of inequalities in (1.7) is estremely useful. Notice that the geometric meaning of (1.7) is that if we restrict our attention to the segment $[x,y]$ with intrinsic real coordinate such that $x\leqslant y$, then $\overline{p}\leqslant s\leqslant\overline{q}$, where $\overline{p}$, $\overline{q}$ are the slopes of the supporting hyperplane at $x,y$ identified by $p,q$ and $s$ is the slope of the chord connecting $(x,u(x))$ and $(y,u(y))$. An immediate consequence of (1.7) is the following monotonicity property.
@@
Since the chain of inequalities (1.7) directly follows from the subdifferentiability of Definition 1.5, Proposition 1.11 tells us that the subdifferential $\partial u$ of any function $u\colon X\to{\mathbb{R}}$ is monotone in the sense that (1.10) holds whenever $x,y$ are such that $\partial u(x)\neq\emptyset\neq\partial u(y)$. In general, this is of no use, since $\partial u$ can be empty in “too many” (possibly all) points. However, if $u$ is convex on $X$ open and convex then Lemma 1.10 shows that the inequality (1.10) is meaningful for each pair of points $x,y\in X$, since the domain of $\partial u$ is the whole $X$; that is, $\mathrm{dom}\,\partial u\vcentcolon=\{x\in X:\partial u(x)\neq\emptyset\}=X$. Furthermore, for differentiable functions $u\colon{\mathbb{R}}\to{\mathbb{R}}$ one knows that convexity is equivalent to the (nondecreasing) monotonicity of $Du=u^{\prime}$, which is in fact equivalent to the monotonicity of $\partial u$ in the sense of (1.10) (cf. also Corollary 1.16 below). Hence, one might wonder whether a characterization of convexity for generic functions exists in term of the monotonicity of their subdifferential. We have essentially just commented on the fact that the answer is no. Indeed, all functions have monotone (possibly empty) subdifferential. Moreover, adding a hypothesis like $\mathrm{dom}\,\partial u\neq\emptyset$ would also be too weak; one needs to assume that $\mathrm{dom}\,\partial u=X$. In fact, consider $u$ defined by $u\vcentcolon=f$ on $X\setminus\{x_{0}\}$ and $u(x_{0})\vcentcolon=\alpha>f(x_{0})$ for some convex function $f$ on $X$. Note that $\mathrm{dom}\,\partial u=X\setminus\{x_{0}\}$ but $u$ is not convex. On the other hand, the assumption $\mathrm{dom}\,\partial u=X$ is too strong because it is in fact itself equivalent to the convexity of $u$. In any case, for the sake of completeness, we mention that a characterization of convexity in terms of the monotonicity of “a subdifferential” actually exists, and uses the so-called proximal subdifferential (see [12, Theorem 4.1] and the references therein).
@@
The implication ($\Longrightarrow$) immediately follows from the monotonicity of the subdifferential (Proposition 1.11) and the upcoming Corollary 1.16 that characterizes convex and differentiable functions. Alternatively, a direct proof is also instructive. The convexity of $u$ is equivalent to the convexity of $u$ restricted to every segment $[x,y]\subset X$ convex with arbitrary endpoints $x,y\in X$; that is, to the convexity of all functions of the form

### v3

@@ Introduction
This is the idea of the monotonicity-duality method initiated by Harvey and Lawson in [26] in the constant coefficient pure second order case, which has developed into a robust and general method which can also be transported to every nonlinear degenerate elliptic PDE which is compatible with a given potential theory. The reader is invited to consult [9, 10, 11, 12, 27] for many of these developments.
We will derive this method from first principles and then discuss the constant coefficient pure second order and gradient-free cases, which also includes the characterization of the dual cone subharmonics, which are the subaffine and subaffine-plus functions, respectively.
The second setting, presented in Section 9.3, treats the opposite situation in which there is only the minimal monotonicity that every subequation possesses. We will prove a weaker form of the comparison principle, called strict comparison. In Theorem 9.19, we treat the variable coefficient case but require that the subharmonics are semiconvex. The proof of strict comparison in this case is a simple consequence of the Summand Theorem 2.16. On the other hand, in the constant coefficient case strict comparison also holds for semicontinuous subharmonics, as shown in Theorem 9.21. The two approaches to strict comparison both use semiconvex approximations; one implicitly and one explicitly.
Part II concludes in Chapter 10 with the final stage which shows why the semiconvex analysis presented in the book extends from Euclidian spaces to manifolds. The key point is Theorem 10.1 which proves that the class of locally semiconvex functions on an open subsets of ${\mathbb{R}}^{n}$ is stable under pull-backs via smooth changes of coordinates.
Commentary
We conclude this introduction with some additional commentary of how this book fits into the rapidly developing and fruitful interplay between general potential theories and the operator theory of fully nonlinear elliptic PDEs. In particular, we will comment on and contrast the parallel developments in general potential theories and the conventional viscosity theory for fully nonlinear elliptic PDEs, including additional comments about some missed opportunities of synergy between the two realms. This will lead to a brief discussion about what is “new” in this primer, with respect to the two unpublished manuscripts [28] and [29]. Finally, we will give a somewhat more detailed history of the progress on the interplay between general potential theories and fully nonlinear elliptic PDEs.
As described above, the “hard analysis part” in treating subharmonics in general potential theories is contained in two deep results on semiconvex functions; namely Alexandrov’s Theorem and the Jensen–Słodkowski Theorem. These two theorems yield the Summand Theorem 2.16 and the Almost Everywhere Theorem 8.1, which ensure that, for locally semiconvex functions, almost all (Alexandrov) points admit an upper test function, and they are enough for the viscosity-theoretic purposes. Hence, locally semiconvex functions can be seen as a potential bridge between the classical theory (twice differentiable functions) and the viscosity theory (semicontinuous functions).
This bridge is completed by the well-known device of approximating a given upper semicontinuous function by a sequence of semiconvex functions by way of the sup-convolution, provided that one has suitable “stability” or “control” of the approximation procedure in order to close the circle and obtain results for the semicontinuous functions. This is a crucial problem, which we will address only in a few basic situations, in Part II. For further investigations on the possibility of applying such a semiconvex approximation technique inside the context of a general potential theory, we invite the reader to consult  [10, 11, 12].
Of course, the conventional viscosity theory for ellipitc PDEs, as codified in [14], also makes use of the same ingredients: Alexandrov’s Theorem, Jensen’s Lemma and “magical properties” of the sup-convolution. However, the more detailed semiconvex analysis presented here can be exploited in a more efficient manner than the conventional approach, as well as providing a more regular “test case” for checking whether a desired result for semicontinuous function can be true. This is accomplished by first proving a (locally) semiconvex version with the help of Alexandrov’s Theorem and the Almost Everywhere Theorem, and then by analyzing the sup-convolution approximation as a last step. In particular, for proving the comparison principle, one can make good use of the structure of the constraint set of a given potential theory (or structural properties of a differential operator) which need not satisfy the standard structural conditions of the conventional viscosity theory. It should be stressed that in the conventional approach to the comparison principle the most delicate step appears in the Theorem on Sums. This is a deep, important and useful general result about the local extrema of sums of upper semicontinuous functions, but in this step the structure of the PDE (or associated potential theory) is completely ignored.
Having recalled and contrasted some of the parallel developments in the viscosity theory of general potential theories and fully nonlinear elliptic PDEs, we repeat our point of view is that an increasing synergy is possible and desirable between the two realms. Historically, there are instances of lost opportunities of such a synergy. The example we will discuss at length concerns the crucial Jensen and Słodkowski Lemmas. Jensen’s Lemma was developed in the context of proving uniqueness for viscosity solutions of fully nonlinear elliptic PDEs [39], while the earlied result of Słodkowski’s Lemma [48] was developed in the context of pluripotential theory. A major aim of this book is to contribute to the many opportunities for productive interplay between these two realms.
@@
We now will comment on what is is “new” in this primer, with respect to the two unpublished manuscripts [28] and [29]. In a general sense, almost everything one finds in the two preprints of Harvey and Lawson is included here, albeit sometimes arranged in a different manner, with modified, clarified and more detailed proofs. Many additional results and remarks have been included to enrich the presentation, which also takes into account the numerous developments in general nonlinear potential theory over the past decade.
@@
In a more technical sense, they are many things here that are not included in [28] and [29]. Some of these are well known, such as elements of convex function theory and the proof of the Lipschitz version of Sard’s theorem, which have been included in order to make the primer as self-contained as reasonably possible. Others are less well known, but needed for the exposition, such as the background material on subequations and their subharmonics in Chapters 5, 6 and 7. Finally, there are some genuinely “new” things, many of which revolve around the following consideration.
@@
A key point in Harvey and Lawson’s study of Jensen–Słodkowski equivalences was to replace the spherical contact points considered by Słodkowski with paraboloidal contact points which are more natural for the viscosity theory. This “paraboloidal” machinery is examined in detail in Section 3.2. In addition, using this machinery, we augment the results in [28]. In particular, we formulate and prove a paraboloidal version of Słodkowski’s upper bound on the largest eigenvalue of the Hessian (Lemma 3.20) and we give a paraboloidal proof of Słodkowski’s density estimate (Lemma 3.2) and Słodkowski’s Largest Eigenvalue Theorem 3.5. Finally, in Section 3.6, making use of an important observation of Harvey [23], we prove that the general Jensen–Słodkowski Theorem 3.27 can be derived by Alexandrov’s maximum principle (Theorem 3.35), which is turn is derived from the Area formula for gradients of semiconvex functions (Theorem 3.34). This leads to the interesting conclusion that, in the end, the hard analysis part rests on the validity of the area formula.
@@
In Part II, an organic presentation of the semiconvex approximation of $\mathcal{F}$ subharmonics is discussed in Chapter 8, which is then used to give an alternate proof of strict comparison for semicontinuous subharmonics (Theorem 9.21) by reducing to the semiconvex version of strict comparison (Theorem 9.19).
We conclude our commentary with a brief history of the program of using general potential theories to treat fully nonlinear elliptic PDEs. The program initiated by Harvey and Lawson was in part inspired by the geometric approach of Krylov [40] in order to formulate a general notion of ellipticity. In the pure second order case, Krylov calls a differential operator $u\mapsto F(D^{2}u)$ elliptic if there exists an elliptic set for $F$; that is, an open subset $\Theta\subset\mathcal{S}(n)$ such that $F$ is increasing for $A\in\Theta$ and the differential inclusion $D^{2}u(x)\in\partial\Theta$ for each $x\in\Omega$ is called an elliptic branch of the equation $F(D^{2}u)=0$ if $\partial\Theta\subset\{F=0\}$. One can then “forget” about the particular form of the operator $F$ and concentrate instead on the constraint set $\Theta$ as many operators $F$ will give rise to the same $\Theta$. Harvey and Lawson recognized similarities of this idea in pluripotential theory and starting instead from closed sets $\mathcal{F}\subset\mathcal{S}(n)$ (the closure of $\Theta$), formulated a suitable notion of duality which recasts the notion of viscosity solutions in a precise topological framework and developed a monotonicity-duality method for proving the comparison principle in general potential theories. With additional work, this yields the comparison principle for any fully nonlinear second-order partial differential equation which is compatible with a given potential theory. The comparison principle of course implies uniqueness of solutions for the associated Dirichlet problem in both contexts (potential theory and PDEs). Another major breakthrough in the Harvey-Lawson approach is to use asymptotic interiors of the constraint set to determine the correct notion of boundary pseudoconvexity for which one can construct the barriers needed to complete Perron’s method for proving existence of solutions for the Dirichlet problem.
After their pioneering paper [26] for pure second order constant coefficient potential theories, a sequence of ambitious papers followed, such as [27] which studies the Dirichlet problem for fully nonlinear second-order equations on Riemannian manifolds, [30] which characterizes the validity of the strong maximum principle, [31] and [34] which treat the special Lagrangian potential equation, [24] on potential theory in calibrated geometry and [32] on the inhomogeneous Dirichlet problem for “natural” operators.
However, the work of Harvey and Lawson (almost always) remains at the level of potential theory, where differential operators appear as possible generators of interesting potential theories and the operators are interesting in and of themselves when they are smooth, so that one can differentiate the differential equation. An important general question concerning the interplay between potential theory and operator theory remained; namely, how to take a given partial differential operator of interest and embed it into a potential theory. Recently, powerful versions of what we call the correspondence principle have been established. The validity of a correspondence principle involves showing that the viscosity subharmonics/superharmonics of a potential theory determined by a subequation $\mathcal{F}$ correspond to the viscosity subsolutions/supersolutions of a differential operator $F$. A correspondence principle is used to embed a given (elliptic) partial differential operator into an associated potential theory and then to deduce properties of the operator theory from the corresponding potential theory. Moreover, these properties are shared by any operator which is compatible with the associated potential theory. Such correspondence principles have been established in the presence of sufficient monotonicity for compatible operator subequation pairs $(F,\mathcal{F})$ in the constant coefficient case in the monograph [9] and for variable coefficient pairs $(F,\mathcal{F})$ under the additional assumption of fiberegularity in the paper [12]. For subequations $\mathcal{F}$ on an open subset $X\subset{\mathbb{R}}^{n}$, fiberegularity just means continuity of the fiber map $\Theta:X\to\mathcal{K}(X)$ defined by
$\Theta(x)\vcentcolon=\mathcal{F}_{x}=\{J\in\mathcal{J}^{2}:(x,J)\in\mathcal{F}\},$
where the closed subsets $\mathcal{K}(X)\subset\mathcal{J}^{2}(X)$ are equipped with the Hausdorff metric.
Inspired by [26], Marco Cirant and the first author began an investigation on embedding large classes of differential operators into the potential theoretic framework of Harvey and Lawson. The variable coefficient pure second order and gradient-free cases were treated in [10] and [11] respectively, where the continuity of the fiber map $x\mapsto\mathcal{F}_{x}=\Theta(x)$ (now called fiberegularity) was shown to be a sufficient additional condition in order to transport the monotonicity-duality method from the constant coefficient to the variable coefficient setting. In addition, a rudimentary version of the correspondence principle was established in [10] and [11] but followed more explicitly the ideas of Krylov ((proper) elliptic maps $\Theta$ and associated branches of the PDE) as opposed to the subequation formulation. Joining forces in [9], a complete and robust treatment of constant coefficient potential theories and PDEs in Euclidian spaces by the monotonicity-duality method was given, including the aforementioned correspondence principle, a classification of monotonicity cone subequations and numerous illustrations of the method. Finally, the general variable coefficient case in the presence of sufficient monotonicity (proper ellipticity and directionality) and fiberegularity is treated by the authors and Cirant in [12].
Part I Semiconvex apparatus
This first part is dedicated to four fundamental and complementary aspects of semiconvex analysis for the viscosity theory of subharmonic functions in general potential theories. The first aspect concerns differentiability properties of first and second order for locally semiconvex functions, including the essential theorem of Alexandrov on the almost everywhere differentiability of second order for (locally semi-)convex functions. The second aspect concerns a detailed analysis of their upper contact jets and upper contact points, culminating in the Upper Contact Jet Theorem and the Summand Theorem. The third aspect concerns the deep analytical results which ensure the existence of sets of contact points with positive measure with the presentation of the Jensen–Słodkowski Theorem and the final aspect presents the well-known device of semiconvex approximation of upper semicontinuous functions, which completes the bridge between the classical and viscosity notions of subharmonics in any potential theory.
Chapter 1 Differentiability of convex functions
In this chapter, we present the well-established theory of differentiability of first and second order for convex functions and their extensions to locally semiconvex functions. This will culminate in the essential theorem of Alexandrov on second order differentiability in the Peano sense on sets of full Lebesgue measure.
1.1. First order theory
We begin by recalling the notion of a convex function and showing that they are subdifferentiable everywhere, where subdifferentiability is a weak and, in general, multi-valued notion. Here we follow and expand upon the treatment in [28]. The results that we are going to discuss actually hold in a more general context, but we limit ourselves to the context that best suits our purposes (see Remark 1.25 below).
Definition 1.1.
Let $u\colon X\to{\mathbb{R}}$ be a function defined on a convex subset $X\subset{\mathbb{R}}^{n}$. We say that $u$ is convex if its epigraph $\operatorname{epi}(u)\vcentcolon=\{y\in{\mathbb{R}}:y\geqslant u(x)\}$ is convex.
We will also say that $u$ is concave if $-u$ is convex.
Remark 1.2 (Reduction to the one-dimensional notion).
It is immediate to see that convexity of $u$ on $X\subset{\mathbb{R}}^{n}$ convex is equivalent to the convexity of the restriction of $u$ to every segment
$[x_{1},x_{2}]\vcentcolon=\{x\in{\mathbb{R}}^{n}:\,x=tx_{1}+(1-t)x_{2},\ t\in[0,1]\},\quad x_{1},x_{2}\in X.$
This translates into the standard inequality formulation for convexity; that is, $u$ is convex in the sense of Definition 1.1 if and only if for each $x_{1},x_{2}\in X$ one has
$u(tx_{1}+(1-t)x_{2})\leqslant tu(x_{1})+(1-t)u(x_{2})\quad\text{for each}\ t\in[0,1].$ (1.1)
Moreover, the inequality (1.1) is a version of Jensen’s inequality for convex functions.
@@ 1.1. First order theory
$u\biggl{(}\,\sum_{j=1}^{N}t_{j}x_{j}\!\biggr{)}\leqslant\sum_{j=1}^{N}t_{j}u(x_{j}),\quad\text{for each}\ \{t_{j}\}_{j=1}^{N}\subset[0,1]\ \text{with}\ \sum_{j=1}^{N}t_{j}=1.$ (1.2)
The proof of Lemma 1.3 is standard and typically done by induction on $N\geqslant 2$, where the case $N=2$ is (1.1).
We also note that the notion of convexity can be localized.
Definition 1.4.
A function $u\colon X\to{\mathbb{R}}$ on an open subset $X\subset{\mathbb{R}}^{n}$ is said to be locally convex if for every $x\in X$ there exists a ball $B\subset X$ centered in $x$ on which $u$ is convex.
Lemma 1.5 (Localization).
Let $u\colon X\to{\mathbb{R}}$ be a function defined on $X$ open and convex. Then
@@
The forward implication is obvious and the reverse implication can be proved by exploiting the restriction of $u$ to arbitrary segments $[x_{1},x_{2}]\subset X$ of Remark 1.2.
We recall now a few fundamental examples of convex functions; examples which can be localized in light of Lemma 1.5
Remark 1.6 (Examples of convex functions).
One easily verifies the following facts:
  1. (a)
    a quadratic function $u$ is convex if and only if its Hessian $D^{2}u$ is a non-negative definite matrix; in particular, every affine function is convex;
  2. (b)
    conical combinations (weighted sums with non-negative coefficients) of a finite family of convex functions are convex;
  3. (c)
    the (pointwise) supremum of any family of convex functions is convex.
@@
Geometrically, this means that the hyperplane which is the graph of the affine function $u(x)+\langle p,\cdot-x\rangle$ over $X$ is a supporting hyperplane from below at $x$ for the epigraph of $u$, for all $p\in\partial u(x)$. One can also say that if $u$ is subdifferentiable at $x$, then $x$ belongs to the (global) lower contact set of $u$, using a terminology which will be introduced in Chapter 2. Moreover, one should note the simple but useful fact that $x$ is a global minimum point for $u$ if and only if $0\in\partial u(x)$. In particular, all functions are subdifferentiable at global minima.
@@
$\partial u\colon X\to\mathscr{P}({\mathbb{R}}^{n}),\quad x\mapsto\partial u(x),$
@@
$u\ \text{is convex}\ \ \Leftrightarrow\ \ \partial u(x)\neq\emptyset\ \ \forall\,x\in X;$
@@
If $u$ is convex, then by the Hahn–Banach Theorem for any $x\in X$ there exists a supporting hyperplane for the convex set $C\vcentcolon=\operatorname{epi}(u)$ at the point $(x,u(x))$; that is, there exists $(q,r)\in{\mathbb{R}}^{n}\times{\mathbb{R}}\setminus\{(0,0)\}$ such that
@@
which yields $-r^{-1}q\in\partial u(x)$, provided that $r\neq 0$. To show that in fact $r\neq 0$, notice that otherwise by inequality (1.4) one would have $\langle q,y-x\rangle\geqslant 0$ for each $y\in X$, which is impossible since $X$ is open.
Conversely, if $x=\lambda x_{1}+(1-\lambda)x_{2}$ for some $\lambda\in[0,1]$, we have
@@
$u$ is bounded on every compact subset $K\subset X$,
@@
$u$ is bounded on some neighborhood of every $x\in X$.
@@
Let $K\subset X$ be compact. By choosing any $x\in K$ and any $p\in\partial u(x)\neq\emptyset$, one has
$u(y)\geqslant u(x)+\langle p,y-x\rangle\vcentcolon=a_{x}(y),\ \ \text{for every}\ y\in X,$
where $a_{x}$ is affine and hence bounded from below on $K$ compact. Hence $u$ is bounded from below on $K$.
For the local boundedness from above, let $x\in X$ be arbitrary and consider any compact cube $\mathcal{C}$ centered in $x$ such that $\mathcal{C}\subset X$. We denote by $\{v_{j}\}_{j=1}^{N}$ the collection of vertices of $\mathcal{C}$. Every closed cube is the (closed) convex hull of its vertices and hence for each $y\in\mathcal{C}$ one has
@@
$u(y)=u\biggl{(}\,\sum_{j=1}^{N}t_{j}v_{j}\biggr{)}\leqslant\sum_{j=1}^{N}t_{j}u(v_{j})\leqslant\max_{1\leqslant j\leqslant N}\lvert u(v_{j})\rvert<+\infty,$
@@
    the fiber $\partial u(x)\subset{\mathbb{R}}^{n}$ is non-empty, closed and convex for each $x\in X$;
@@
$u(y)\geqslant u(x)+\langle p,y-x\rangle\quad\forall p\in\partial u(x),\ \forall\,y\in X$
@@
$u(x)\geqslant u(y)+\langle q,x-y\rangle\quad\forall q\in\partial u(y),\ \forall\,x\in X,$
@@
We will see that the chain of inequalities in (1.5) is extremely useful. Notice that the geometric meaning of (1.5) is that if we restrict our attention to the segment $[x,y]$ with intrinsic real coordinate such that $x\leqslant y$, then $\overline{p}\leqslant s\leqslant\overline{q}$, where $\overline{p}$, $\overline{q}$ are the slopes of the supporting hyperplane at $x,y$ identified by $p,q$ and $s$ is the slope of the chord connecting $(x,u(x))$ and $(y,u(y))$. An immediate consequence of (1.5) is the following monotonicity property.
@@
Since the chain of inequalities (1.5) directly follows from the subdifferentiability of Definition 1.7, Proposition 1.13 tells us that the subdifferential $\partial u$ of any function $u\colon X\to{\mathbb{R}}$ is monotone, in the sense that (1.6) holds whenever $x,y$ are such that $\partial u(x)\neq\emptyset\neq\partial u(y)$. In general, this is of no use, since $\partial u$ can be empty in “too many” (possibly all) points. However, if $u$ is convex on $X$ open and convex then Lemma 1.12 shows that the inequality (1.6) is meaningful for each pair of points $x,y\in X$, since the domain of $\partial u$ is the whole $X$; that is, $\mathrm{dom}\,\partial u\vcentcolon=\{x\in X:\partial u(x)\neq\emptyset\}=X$. Furthermore, for differentiable functions $u\colon{\mathbb{R}}\to{\mathbb{R}}$ one knows that convexity is equivalent to the (nondecreasing) monotonicity of $Du=u^{\prime}$, which is in fact equivalent to the monotonicity of $\partial u$ in the sense of (1.6) (cf. also Corollary 1.18 below). Hence, one might wonder whether a characterization of convexity for generic functions exists in term of the monotonicity of their subdifferential. We have essentially just commented on the fact that the answer is no. Indeed, all functions have monotone (possibly empty) subdifferential. Moreover, adding a hypothesis like $\mathrm{dom}\,\partial u\neq\emptyset$ would also be too weak; one needs to assume that $\mathrm{dom}\,\partial u=X$. In fact, consider $u$ defined by $u\vcentcolon=f$ on $X\setminus\{x_{0}\}$ and $u(x_{0})\vcentcolon=\alpha>f(x_{0})$ for some convex function $f$ on $X$. Note that $\mathrm{dom}\,\partial u=X\setminus\{x_{0}\}$ but $u$ is not convex. On the other hand, the assumption $\mathrm{dom}\,\partial u=X$ is too strong because it is in fact itself equivalent to the convexity of $u$. In any case, for the sake of completeness, we mention that a characterization of convexity in terms of the monotonicity of “a subdifferential” actually exists, and uses the so-called proximal subdifferential (see [13, Theorem 4.1] and the references therein).
@@
The implication ($\Longrightarrow$) immediately follows from the monotonicity of the subdifferential (Proposition 1.13) and the upcoming Corollary 1.18 that characterizes convex and differentiable functions. Alternatively, a direct proof is also instructive: the convexity of $u$ is equivalent to the convexity of $u$ restricted to every segment $[x,y]\subset X$ convex with arbitrary endpoints $x,y\in X$; that is, to the convexity of all functions of the form